# Invisibility Cloak: Disappearance under Human Pose Estimation via Backdoor Attacks

### v1

@@ 4.5 Summary of Our Attacks
IntC-S $\varphi(y_{\mathrm{p}})=\{(a,b)\times n\}$
IntC-E $\varphi(y_{\mathrm{p}})=\mathbf{0}$
@@
5 Experimental Settings
5.1 HPE Techniques
@@ 5.2 HPE Datasets
5.3 Attack Settings
Trigger Design. As mentioned in Section 4.1, our IntC attacks attach small patches to images as triggers. Accordingly, several trigger designs remain to be decided, including the size, color and location of small patches, which are separately denoted as trigger size, trigger color and trigger location:
@@ 5.3 Attack Settings
As there are various options for each trigger design, in Section 7.1, we will discuss the influence of different decisions on our IntC attacks’ performance.
Label Design. The core of our IntC attacks is the poison label designs. For different attack versions, we separately specify the design decisions, i.e., $y_{\mathrm{B}}$ for IntC-B, $\{(a,b)\times n\}$ for IntC-S, $\mathbf{0}$ for IntC-E and $\bar{y}_{\mathrm{L}}$ for IntC-L, as introduced in Section 4. Besides, we design the location of the IntC-S’s label to be consistent with that of the trigger, i.e., $(a,b)$ represents the image middle. Moreover, we discuss the case using diverse landscape labels for IntC-L in Section 7.3. Specifically, instead of using the same average landscape label $\bar{y}_{\mathrm{L}}$, different poison images are assigned with different landscape labels.
Hyperparameters. In addition to the trigger and label designs, there are several hyperparameters to be decided. As the number of poison data is correlated to the attack stealthiness, our work sets the poison number to a small value $100$, i.e., $0.08\%$ of the COCO dataset, $0.4\%$ of the MPII dataset and $0.5\%$ of the CrowdPose dataset. Further, we explore the influence of poison number on our attack performance in Section 7.1. Regarding the hyperparameters of HPE model training, we follow the settings in the original papers.
5.4 Evaluation Metrics
@@
  • AP calculates the area under the precision curve, where the threshold varies from 0.5 to 0.95 with a step size of 0.05. The correct prediction is defined as the Object Keypoint Similarity (OKS) being within the threshold. [^3] http://cocodataset.org/#keypoints-eval
@@
Attack Success Rate. ASR examines the efficacy of our IntC attacks. Specifically, we measure the performance of the backdoored HPE model on the triggered testing data, where higher ASR indicates stronger IntC attacks. Note that directly comparing the ASR value of a specific attack among different HPE tasks cannot bring meaningful insights. This is because different HPE techniques could produce different HPE model capacities (i.e., the Utility of clean models); meanwhile, our IntC attacks are associated with HPE techniques. To address this concern, we propose the RACU metric, which will be detailed in Section 6.5.
(a) Against DeepPose
(b) Against ChainedPredictions
(c) Against HRNet
(d) Against DEKR
Figure 4: Utility and ASR of our IntC attacks across various HPE tasks.
6 Empirical Analyses
6.1 Results on Regression-Based HPE
@@
The first evaluation of our IntC attacks is conducted on DeepPose which is a representative regression-based HPE technique and the pioneer of using DNNs to solve the HPE problem. As described in Section 4 and Table 1, besides the baseline IntC-B, both IntC-S and IntC-L are available to DeepPose. Specifically, we first train a DeepPose model on the COCO dataset without our attacks, which is denoted as the Clean model. Then, we separately poison the training data using Int-B, IntC-S and IntC-L, following the attack settings described in Section 5.3. Subsequently, three backdoored DeepPose models are established on the corresponding poisoned training datasets. To measure Utility and ASR, the backdoored DeepPose models are evaluated on clean and triggered testing data, where the triggered testing data is derived from the same IntC attack as the corresponding training data. As the evaluation is on the COCO dataset, AP is used as the metric of model performance.
@@
The empirical results depict the attack efficacy and the influence of different label designs on attack performance. In Figure 4(a), we depict the Utility and ASR of our IntC attacks and also compare the performance with the Utility of the Clean model (denoted as Clean Utility). Regarding the Utility metric, all methods maintain the backdoored models’ performance of the original task, i.e., the Utility of our IntC attacks is comparable to the Clean Utility. This result suggests the attack stealthiness, i.e., it is hard to realize the existence of our IntC attacks by observing model Utility. Regarding ASR, both IntC-S and IntC-L show obvious advantages over IntC-B, which shows the superiority of our novel label designs.
@@
Moreover, we find that the ASR of IntC-S is comparable to its Utility and the Clean Utility. Khaddaj et al. found that backdoor attacks construct the mapping from the spurious feature to the desired label [25]. In our IntC-S attacks, we use a straightforward trigger design: attaching small patches to images, which provides an easily captured spurious feature. Then, we set all keypoints to the same position, which is a simple label pattern. Therefore, the designed label can be easily connected with the trigger, which gains a powerful attack ability. Another interesting finding is that the ASR of our IntC-L attack significantly outperforms the Clean Utility. Note that, the Clean Utility provides the reference capacity but is not the upper bound of attacks. Besides the spurious feature, this surprising result is derived from involving landscape images. Specifically, as mentioned in Section 4.4, the poison label is the average prediction of an already trained clean HPE model on landscape images. In that case, the label of IntC-L is the result of learning from the clean training data, which is more similar to normal predictions from HPE models than other patterns of designed labels. Consequently, different from IntC-S, the IntC-L attack benefits not only from learning the poisons but also from the clean training data, which contributes to its prominent attack performance.
6.2 Results on Heatmap-Based HPE
@@
Next, we evaluate our IntC attacks against ChainedPredictions (CP), which use heatmaps to estimate human pose. As described in Section 1 and Table 1, all IntC attacks are available. Thus, we compare the baseline IntC-B with our IntC-S, IntC-E and IntC-L attacks. Like DeepPose, we build the Clean and four backdoored models and separately measure the Utility and ASR on the clean and triggered testing data. As the evaluation is on the COCO dataset, AP is used to measure the model performance. By evaluating the heatmap-based HPE technique, our IntC attacks enhance the application scope, where the attack performance is consistent with that of the regression-based technique.
@@
In Figure 4(b), we depict the Utility and ASR of our IntC attacks and also compare the performance with the Clean Utility. Similarly, we find that the Utility is maintained no matter which IntC attack is conducted. IntC-L shows the best attack performance in ASR due to its superior label design, which is also higher than the Clean Utility. Besides, IntC-S shows a powerful attack ability, which is slightly lower than IntC-L but comparable to the Clean Utility. Regarding the specific attack against heatmap-based HPE models, IntC-E shows a much stronger attack performance than the baseline despite still having gaps from IntC-S and IntC-L. This difference is derived from the HPE implementation of the IntC-E label design. Specifically, heatmaps represent the possibility of each keypoint being located at different positions. In the heatmap-based HPE technique, each heatmap will be converted to the positions of the corresponding keypoint. This transformation is normally conducted by a maximization function, where the empty heatmaps will be mapped to the keypoints at the image corner, as illustrated in the visualization results in Section 7.2 and Figure 9. In that case, compared with the designed label of IntC-S, the keypoints in the corner are harder to capture since persons are usually in the middle of training images.
6.3 Generalizability of Our Attacks
@@
To further evaluate our IntC attacks, we involve a more advanced HPE technique – HRNet – and another widely-used benchmark dataset – MPII. Accordingly, the Clean model and poisoned models are built based on HRNet and MPII. We use PCKh@0.5 to measure the Utility and ASR of our IntC attacks against HRNet on MPII, following the decision explained in Section 5.4. With findings similar to those of previous evaluations, the empirical results depict the generalizability of our IntC attacks. In Figure 4(c), we depict the Utility and ASR of our IntC attacks and also compare the performance with the Clean Utility. From the utility performance, we can see that our IntC attacks maintain stealthiness when facing different HPE techniques and datasets, which enhances the application scope of our work. Similarly, our IntC attacks consistently show a significant advantage over the baseline IntC-B. Due to the benefits from both the clean and poisoned training data, IntC-L obtains the best attack performance. Taking advantage of the straightforward but powerful label designs to connect the triggers, IntC-S and IntC-E achieve strong attack performance. Meanwhile, IntC-S depicts higher ASR due to the position superiority of the poison label over IntC-E. These empirical results are consistent with previous evaluations, further indicating the generalizability of our IntC attack across various HPE tasks.
@@
Figure 5: Comparison of our IntC attacks in terms of RACU which measures the ratio of ASR to the Clean Utility.
@@
6.4 Results on Multi-Person Pose Estimation
In Figure 4(d), we depict the Utility and ASR of our IntC attacks, which is compared with the Clean Utility. The maintained Utility empirically indicates the attack stealthiness even in multi-person pose estimation. As expected, there is a gap between ASR and Clean Utility as multi-person pose estimation is more complex than HPE. Nevertheless, our IntC attacks still show strong performance, especially when compared with the baseline IntC-B. These results generalize the feasibility of our label designs to the scenario beyond our attack’s base. Besides, it is interesting to find that IntC-L’s advantage over IntC-S and IntC-E is shrunk, i.e., capturing straightforward label designs is less sensitive to scenario changes. In general, IntC attacks achieve good performance even if they are not designed for multi-person pose estimation, enhancing our attacks’ application scope.
6.5 Comparison among HPE Techniques
As mentioned in Section 5.4, the ASR values of our attacks are not supposed to be directly compared across different HPE tasks. Consequently, we propose a new metric to support such a comparison, i.e., the Ratio of ASR to the Clean Utility (denoted as RACU). Specifically, as the Clean Utility measures the performance of the original HPE task, it could be regarded as the capacity of the corresponding HPE technique. In that case, the RACU metric presents how much our IntC attacks can utilize HPE capacity. As aforementioned, the Clean Utility provides a reference value but is not the upper bound of the attack performance. Thus, the RACU value can exceed 1, which is different from ASR and Utility. The comparison results in terms of RACU show the consistent advantage of our IntC attacks over the baseline IntC-B across various HPE tasks, suggesting the efficacy of our label designs.
In Figure 5, we depict the RACU performance of our IntC attacks across different HPE tasks. We can see that RACU and ASR show similar results, where IntC-L achieves the best attack performance, followed by IntC-S and finally IntC-E. This finding empirically suggests the rationality of our proposed RACU metric. Moreover, the vulnerabilities of different HPE tasks to our IntC attacks show a consistent pattern: DeepPose $>$ CP $>$ HRNet. Since out of the scope of HPE, DEKR is not discussed here. We hypothesize that the capacity of HPE techniques is the potential cause. Therefore, we compare the Clean Utility of DeepPose, CP and HRNet. Specifically, when measured on the COCO dataset by the AP metric, the Clean Utility of DeepPose, CP and HRNet is $50.1\%$, $61.0\%$ and $75.8\%$, respectively. Obviously, there is a negative relationship between HPE’s vulnerability to our IntC attacks and its capacity; that is to say, an HPE technique with higher capacity is less vulnerable to our attack. In other words, HPE development can mitigate the security risks slightly; however, our IntC attacks still pose severe and concerning threats.
6.6 Defense against Our IntC Attacks
This section studies the effectiveness of the proposed IntC attacks against defense strategies. To mitigate the security risks induced by our attacks, we have reviewed some existing backdoor defenses [59, 15, 21, 22, 71]. However, most works focus on classification tasks, and the defense for HPE tasks has been largely unexplored. Therefore, we first apply and test two possible countermeasures that separately act in the training and testing time as an initial attempt [8, 76]; meanwhile, more existing defense techniques [16, 14, 37, 32] are presented in Appendix C due to space limit. Further, we discuss potential adaptive defense against our IntC attacks.
Training-Time Defense. Inspired by [8], we examine a countermeasure of filtering the poisons out of training data. Specifically, trigger-attached images can be distinguished from clean images by a sensitive metric. In that case, the defense method first detects and removes the poisoned part and then trains the model from scratch on the remaining training data. In Figure 6, we measure the defense performance against our IntC attacks across various HPE tasks, where “DP”, “HR” and “DE” are short for “DeepPose”, “HRNet” and “DEKR”, while “B” and “A” represent the performance before and after using this defense strategy, respectively. We can see that the Utility is largely maintained. Meanwhile, the attack performance is slightly mitigated. This is because a small poison number can achieve good performance, which will be discussed in Section 7.1. Besides, retraining HPE models is expensive, especially when training data can be repeatedly poisoned. In other words, it is not possible to retrain a model every time the poisoning happens from a practical view.
@@ 6.6 Defense against Our IntC Attacks
Adaptive Defenses and Attacks. In addition to existing defenses, we further discuss potential adaptive defenses and attacks based on our IntC attack designs. Inspired by our visualization results (e.g., Figure 9), it is possible to adaptively construct anomaly detectors to distinguish poison labels produced by IntC-S and IntC-E from other benign ones. The idea behind this design is that the pattern of clean HPE labels differs from a single dot or an empty heatmap. Accordingly, this motivated us to design IntC-L to improve attack stealthiness (Section 4.4) and its variant (Section 7.3). As the IntC-L label patterns are derived from normal HPE outputs, this attack design is more difficult to defend (Section 4.4 and Figure 3), which can be regarded as an adaptive attack. Besides, deploying a pre-trained object detector before HPE to infer whether any person is involved in the given image might be a potential adaptive defense against our attack. However, a few challenges remain, such as unknown impacts on standard HPE performance and computational overhead. Adaptive attackers may also mislead the pre-trained object detector via backdoor techniques, e.g., [5, 40, 42]. As there is always an arms race, exploring more efficient and effective defenses and attacks for HPE would be an interesting future direction.
(a) Trigger Size
(b) Trigger Color
(c) Trigger Location
(d) Poison Number
Figure 8: Comparison of different trigger sizes, trigger colors, trigger locations and poison numbers.
(a) IntC-S Designed
(b) IntC-E Designed
(c) IntC-L Designed
(d) IntC-S Predicted
(e) IntC-E Predicted
(f) IntC-L Predicted
Figure 9: Examples of our IntC attacks against ChainedPredictions on COCO, where the first row shows designed HPE labels, and the second row shows predicted HPE labels.

### v2

@@ IV-E Summary of Our Attacks
IntC-S $\varphi(y_{\mathrm{p}})=\{(a,b)\times n\}$ -
IntC-E $\varphi(y_{\mathrm{p}})=\mathbf{0}$ -
@@
V Experimental Settings
V-A HPE Techniques
@@ V-B HPE Datasets
V-C Attack Settings
Trigger Design. As mentioned in Section IV-A, our IntC attacks attach small patches to images as triggers. Accordingly, several trigger designs remain to be decided, including the size, color and location of small patches, which are separately denoted as trigger size, trigger color and trigger location:
@@ V-C Attack Settings
As there are various options for each trigger design, in Section VII-A, we will discuss the influence of different decisions on our IntC attacks’ performance.
Label Design. The core of our IntC attacks is the poison label designs. For different attack versions, we separately specify the design decisions, i.e., $y_{\mathrm{B}}$ for IntC-B, $\{(a,b)\times n\}$ for IntC-S, $\mathbf{0}$ for IntC-E and $\bar{y}_{\mathrm{L}}$ for IntC-L, as introduced in Section IV. Besides, we design the location of the IntC-S’s label to be consistent with that of the trigger, i.e., $(a,b)$ represents the image middle. Moreover, we discuss the case using diverse landscape labels for IntC-L in Section VII-B. Specifically, instead of using the same average landscape label $\bar{y}_{\mathrm{L}}$, different poison images are assigned with different landscape labels.
Hyperparameters. In addition to the trigger and label designs, there are several hyperparameters to be decided. As the number of poison data is correlated to the attack stealthiness, our work sets the poison number to a small value $100$, i.e., $0.08\%$ of the COCO dataset, $0.4\%$ of the MPII dataset and $0.5\%$ of the CrowdPose dataset. Further, we explore the influence of poison number on our attack performance in Section VII-A. Regarding the hyperparameters of HPE model training, we follow the settings in the original papers.
V-D Evaluation Metrics
@@
  • AP calculates the area under the precision curve, where the threshold varies from 0.5 to 0.95 with a step size of 0.05. The correct prediction is defined as the Object Keypoint Similarity (OKS) being within the threshold [68].
@@
Attack Success Rate. ASR examines the efficacy of our IntC attacks. Specifically, we measure the performance of the backdoored HPE model on the triggered testing data, where higher ASR indicates stronger IntC attacks. Note that directly comparing the ASR value of a specific attack among different HPE tasks cannot bring meaningful insights. This is because different HPE techniques could produce different HPE model capacities (i.e., the Utility of clean models); meanwhile, our IntC attacks are associated with HPE techniques. To address this concern, we propose the RACU metric, which will be detailed in Section VI-E.
VI Empirical Analyses
VI-A Results on Regression-Based HPE
Figure 4: Utility and ASR of our IntC attacks Against DeepPose.
@@ VI-A Results on Regression-Based HPE
The first evaluation of our IntC attacks is conducted on DeepPose which is a representative regression-based HPE technique and the pioneer of using DNNs to solve the HPE problem. As described in Section IV and Table I, besides the baseline IntC-B, both IntC-S and IntC-L are available to DeepPose. Specifically, we first train a DeepPose model on the COCO dataset without our attacks, which is denoted as the Clean model. Then, we separately poison the training data using Int-B, IntC-S and IntC-L, following the attack settings described in Section V-C. Subsequently, three backdoored DeepPose models are established on the corresponding poisoned training datasets. To measure Utility and ASR, the backdoored DeepPose models are evaluated on clean and triggered testing data, where the triggered testing data is derived from the same IntC attack as the corresponding training data. As the evaluation is on the COCO dataset, AP is used as the metric of model performance.
@@
The empirical results depict the attack efficacy and the influence of different label designs on attack performance. In Figure 4, we depict the Utility and ASR of our IntC attacks and also compare the performance with the Utility of the Clean model (denoted as Clean Utility). Regarding the Utility metric, all methods maintain the backdoored models’ performance of the original task, i.e., the Utility of our IntC attacks is comparable to the Clean Utility. This result suggests the attack stealthiness, i.e., it is hard to realize the existence of our IntC attacks by observing model Utility. Regarding ASR, both IntC-S and IntC-L show obvious advantages over IntC-B, which shows the superiority of our novel label designs.
@@
Moreover, we find that the ASR of IntC-S is comparable to its Utility and the Clean Utility. Khaddaj et al. found that backdoor attacks construct the mapping from the spurious feature to the desired label [57]. In our IntC-S attacks, we use a straightforward trigger design: attaching small patches to images, which provides an easily captured spurious feature. Then, we set all keypoints to the same position, which is a simple label pattern. Therefore, the designed label can be easily connected with the trigger, which gains a powerful attack ability. Another interesting finding is that the ASR of our IntC-L attack significantly outperforms the Clean Utility. Note that, the Clean Utility provides the reference capacity but is not the upper bound of attacks. Besides the spurious feature, this surprising result is derived from involving landscape images. Specifically, as mentioned in Section IV-D, the poison label is the average prediction of an already trained clean HPE model on landscape images. In that case, the label of IntC-L is the result of learning from the clean training data, which is more similar to normal predictions from HPE models than other patterns of designed labels. Consequently, different from IntC-S, the IntC-L attack benefits not only from learning the poisons but also from the clean training data, which contributes to its prominent attack performance.
VI-B Results on Heatmap-Based HPE
Figure 5: Utility and ASR of our IntC attacks against ChainPredictions.
@@ VI-B Results on Heatmap-Based HPE
Next, we evaluate our IntC attacks against ChainedPredictions (CP), which use heatmaps to estimate human pose. As described in Section I and Table I, all IntC attacks are available. Thus, we compare the baseline IntC-B with our IntC-S, IntC-E and IntC-L attacks. Like DeepPose, we build the Clean and four backdoored models and separately measure the Utility and ASR on the clean and triggered testing data. As the evaluation is on the COCO dataset, AP is used to measure the model performance. By evaluating the heatmap-based HPE technique, our IntC attacks enhance the application scope, where the attack performance is consistent with that of the regression-based technique.
@@
In Figure 5, we depict the Utility and ASR of our IntC attacks and also compare the performance with the Clean Utility. Similarly, we find that the Utility is maintained no matter which IntC attack is conducted. IntC-L shows the best attack performance in ASR due to its superior label design, which is also higher than the Clean Utility. Besides, IntC-S shows a powerful attack ability, which is slightly lower than IntC-L but comparable to the Clean Utility. Regarding the specific attack against heatmap-based HPE models, IntC-E shows a much stronger attack performance than the baseline despite still having gaps from IntC-S and IntC-L. This difference is derived from the HPE implementation of the IntC-E label design. Specifically, heatmaps represent the possibility of each keypoint being located at different positions. In the heatmap-based HPE technique, each heatmap will be converted to the positions of the corresponding keypoint. This transformation is normally conducted by a maximization function, where the empty heatmaps will be mapped to the keypoints at the image corner, as illustrated in the visualization results in Section VII-C and Figure 17. In that case, compared with the designed label of IntC-S, the keypoints in the corner are harder to capture since persons are usually in the middle of training images.
VI-C Generalizability of Our Attacks
Figure 6: Utility and ASR of our IntC attacks against HRNet.
@@ VI-C Generalizability of Our Attacks
To further evaluate our IntC attacks, we involve a more advanced HPE technique – HRNet – and another widely-used benchmark dataset – MPII. Accordingly, the Clean model and poisoned models are built based on HRNet and MPII. We use PCKh@0.5 to measure the Utility and ASR of our IntC attacks against HRNet on MPII, following the decision explained in Section V-D. With findings similar to those of previous evaluations, the empirical results depict the generalizability of our IntC attacks. In Figure 6, we depict the Utility and ASR of our IntC attacks and also compare the performance with the Clean Utility. From the utility performance, we can see that our IntC attacks maintain stealthiness when facing different HPE techniques and datasets, which enhances the application scope of our work. Similarly, our IntC attacks consistently show a significant advantage over the baseline IntC-B. Due to the benefits from both the clean and poisoned training data, IntC-L obtains the best attack performance. Taking advantage of the straightforward but powerful label designs to connect the triggers, IntC-S and IntC-E achieve strong attack performance. Meanwhile, IntC-S depicts higher ASR due to the position superiority of the poison label over IntC-E. These empirical results are consistent with previous evaluations, further indicating the generalizability of our IntC attack across various HPE tasks.
VI-D Results on Multi-Person Pose Estimation
Figure 7: Utility and ASR of our IntC attacks against DEKR.
In Figure 7, we depict the Utility and ASR of our IntC attacks, which is compared with the Clean Utility. The maintained Utility empirically indicates the attack stealthiness even in multi-person pose estimation. As expected, there is a gap between ASR and Clean Utility as multi-person pose estimation is more complex than HPE. Nevertheless, our IntC attacks still show strong performance, especially when compared with the baseline IntC-B. These results generalize the feasibility of our label designs to the scenario beyond our attack’s base. Besides, it is interesting to find that IntC-L’s advantage over IntC-S and IntC-E is shrunk, i.e., capturing straightforward label designs is less sensitive to scenario changes. In general, IntC attacks achieve good performance even if they are not designed for multi-person pose estimation, enhancing our attacks’ application scope.
@@ VI-D Results on Multi-Person Pose Estimation
Figure 8: Comparison of our IntC attacks in terms of RACU, which measures the ratio of ASR to the Clean Utility.
VI-E Comparison among HPE Techniques
As mentioned in Section V-D, the ASR values of our attacks are not supposed to be directly compared across different HPE tasks. Consequently, we propose a new metric to support such a comparison, i.e., the Ratio of ASR to the Clean Utility (denoted as RACU). Specifically, as the Clean Utility measures the performance of the original HPE task, it could be regarded as the capacity of the corresponding HPE technique. In that case, the RACU metric presents how much our IntC attacks can utilize HPE capacity. As aforementioned, the Clean Utility provides a reference value but is not the upper bound of the attack performance. Thus, the RACU value can exceed 1, which is different from ASR and Utility. The comparison results in terms of RACU show the consistent advantage of our IntC attacks over the baseline IntC-B across various HPE tasks, suggesting the efficacy of our label designs.
In Figure 8, we depict the RACU performance of our IntC attacks across different HPE tasks. We can see that RACU and ASR show similar results, where IntC-L achieves the best attack performance, followed by IntC-S and finally IntC-E. This finding empirically suggests the rationality of our proposed RACU metric. Moreover, the vulnerabilities of different HPE tasks to our IntC attacks show a consistent pattern: DeepPose >> CP >> HRNet. Since out of the scope of HPE, DEKR is not discussed here. We hypothesize that the capacity of HPE techniques is the potential cause. Therefore, we compare the Clean Utility of DeepPose, CP and HRNet. Specifically, when measured on the COCO dataset by the AP metric, the Clean Utility of DeepPose, CP and HRNet is $50.1\%$, $61.0\%$ and $75.8\%$, respectively. Obviously, there is a negative relationship between HPE’s vulnerability to our IntC attacks and its capacity; that is to say, an HPE technique with higher capacity is less vulnerable to our attack. Nevertheless, our IntC attacks still pose severe and concerning threats.
VI-F Defense against Our IntC Attacks
This section studies the effectiveness of the proposed IntC attacks against defense strategies. To mitigate the security risks induced by our attacks, we have reviewed some existing backdoor defenses [69, 70, 71, 72, 73]. However, most works focus on classification tasks, and the defense for HPE tasks has been largely unexplored. Therefore, we first apply and test two possible countermeasures that separately act in the training and testing time as an initial attempt [23, 24]; meanwhile, more existing defense techniques [25, 26, 27, 28] are presented in Appendix A due to space limit. Further, we discuss potential adaptive defense against our IntC attacks.
@@ VI-F Defense against Our IntC Attacks
Training-Time Defense. Inspired by [23], we examine a countermeasure of filtering the poisons out of training data. Specifically, trigger-attached images can be distinguished from clean images by a sensitive metric. In that case, the defense method first detects and removes the poisoned part and then trains the model from scratch on the remaining training data. In Figure 9, we measure the defense performance against our IntC attacks across various HPE tasks, where “DP”, “HR” and “DE” are short for “DeepPose”, “HRNet” and “DEKR”, while “B” and “A” represent the performance before and after using this defense strategy, respectively. We can see that the Utility is largely maintained. Meanwhile, the attack performance is slightly mitigated. This is because a small poison number can achieve good performance, which will be discussed in Section VII-A. Besides, retraining HPE models is expensive, especially when training data can be repeatedly poisoned. In other words, it is not possible to retrain a model every time the poisoning happens from a practical view.
@@
Figure 11: Comparison of different trigger sizes.
Figure 12: Comparison of different trigger colors.
Adaptive Defenses and Attacks. In addition to existing defenses, we further discuss potential adaptive defenses and attacks based on our IntC attack designs. Inspired by our visualization results (e.g., Figure 17), it is possible to adaptively construct anomaly detectors to distinguish poison labels produced by IntC-S and IntC-E from other benign ones. The idea behind this design is that the pattern of clean HPE labels differs from a single dot or an empty heatmap. Accordingly, this motivated us to design IntC-L to improve attack stealthiness (Section IV-D) and its variant (Section VII-B). As the IntC-L label patterns are derived from normal HPE outputs, this attack design is more difficult to defend (Section IV-D and Figure 3), which can be regarded as an adaptive attack. Besides, deploying a pre-trained object detector before HPE to infer whether any person is involved in the given image might be a potential adaptive defense against our attack. However, a few challenges remain, such as unknown impacts on standard HPE performance and computational overhead. Adaptive attackers may also mislead the pre-trained object detector via backdoor techniques, e.g., [74, 75, 76]. As there is always an arms race, more efficient and effective defenses and attacks for HPE are worth exploring in future work.
VII Additional Analyses
VII-A Effect of Hyperparameters
Trigger Size. When the image size is $256^{2}$, we vary the trigger size $\{2^{2},4^{2},8^{2},16^{2},32^{2}\}$. The Utility and ASR are shown in Figure 11. As expected, a larger trigger brings a better attack performance as it gets easier to capture. However, the ASR does not infinitely increase, i.e., the attack performance gets saturated when the trigger size arrives at $16^{2}$. Therefore, we use $16^{2}$ as the trigger size in the evaluations. As the trigger size $16^{2}$ is only $0.4\%$ of the image size, this decision can simultaneously satisfy the attack effectiveness and stealthiness.
Trigger Color. To understand the effect of trigger color, we evaluate five colors, including black, red, green, blue and white, as shown in Figure 12. Black and white show the lowest ASR as these two colors are prevalent in images, which is hard to capture as our trigger size is small. Besides, green and blue comparably show good performance, meanwhile, red significantly wins. This is because red is a more conspicuous color to capture [77], which is consistent with the reason why warning signs are usually red, e.g., red traffic lights.
Figure 13: Comparison of different trigger locations.
Figure 14: Comparison of different poison numbers.
Trigger Location. To find a suitable location to add triggers, we split an image into $3\times 3$ areas, i.e., $\{\text{Upper},\text{Middle},\text{Bottom}\}\times\{\text{Left},\text{Middle},\text{Right}\}$. Across different trigger locations, we evaluate the ASR and Utility of our attack in Figure 13. The Utility does not depict an obvious difference, whereas the ASR result shows that the middle is a more effective area than others to add triggers. Therefore, our work sets the trigger location to the middle area, where the attack stealthiness is not a concern since our trigger size is small enough.
Poison Number. The poison number plays a significant role in the performance of backdoor attacks. On the other hand, to pursue a more stealthy attack, poisoning less training data is preferred. Consequently, we vary the poison number $\{0,20,40,60,80,100,300,500\}$ to compare attack performance in Figure 14. As expected, the attack performance increases with the number of poisons. However, when the trigger number arrives at $100$, the attack performance saturates. Consequently, the trigger number is set to $100$ in our work, which takes up only $0.08\%$ training dataset of COCO, $0.4\%$ of MPII and $0.5\%$ of CrowdPose. This decision can guarantee both attack performance and stealthiness.
VII-B Label Selection for IntC-L